a private airplane flew 420 mi in 3 hours with the wind. The return trip against the wind took 3.5 hours. If the rate of the wind was the same on both trips and the plane flew at the same rate (speed) on both trips, what is the rate (speed) of the wind and the rate (speed) of the plane?
Answer by Alan3354(69443)

a private airplane flew 420 mi in 3 hours with the wind. The return trip against the wind took 3.5 hours. If the rate of the wind was the same on both trips and the plane flew at the same rate (speed) on both trips, what is the rate (speed) of the wind and the rate (speed) of the plane?
----------------
420/3 = 140 mi/hr
420/3.5 = 120 mi/hr
The airspeed of the plane is the average = (140 + 120)/2 = 130 mi/hr
The windspeed is the difference (140 - 130) = 10 mi/hr